# Hint of Thought prompting: an explainable and zero-shot approach to reasoning tasks with LLMs

### v6

@@ 7.3 Discussion about Existing Work
Recently, there have been many approaches to enhance the reasoning ability of LLMs, including CoT Wei et al. (2022a), zero-shot CoT Kojima et al. (2022), Auto-CoT Shin et al. (2020), PoT Chen et al. (2023), decomposed prompting Khot et al. (2023). They all aim to provide accurate reasoning results. On the other hand, they did not focus on answer explanation and the LLM efficiency. Therefore, HoT is proposed and gives another approach to prompt engineering.
@@ 8 Discussion
In our work, we approach the reasoning tasks from other points of view. We care about accuracy and want a more precise explanation of the reasoning process. We have verified that HoT can work efficiently on arithmetic and common sense reasoning tasks and provide a clear description intuitively. Although LLMs are black boxes, we tried to explain them with a prompt approach.
@@ 9 Conclusion
This paper investigates how to generate a more intuitively explainable end-to-end prompting on reasoning tasks. Our experimental results on the mathematical reasoning tasks significantly surpass that of prior zero-shot methods with GPT-3.5 family, CoT, and PoT. Also, on the common sense benchmark StrategyQA, our HoT reaches a new zero-shot GPT-3.5 result with 82.96%.

### v7

@@ 7.3 Discussion about Existing Work
Recently, there have been many approaches to enhance the reasoning ability of LLMs, including CoT Wei et al. (2022), zero-shot CoT Kojima et al. (2022), Auto-CoT Shin et al. (2020), PoT Chen et al. (2023), decomposed prompting Khot et al. (2023). They all aim to provide accurate reasoning results. On the other hand, they did not focus on explanations of answers and the LLM efficiency. Therefore, HoT is proposed as another approach to prompt engineering.
@@ 8 Discussion
In our work, we approach reasoning tasks from other perspectives. We care about accuracy and want a more precise explanation of the reasoning process. We have verified that HoT can work efficiently on arithmetic and common-sense reasoning tasks and provide a clear description intuitively. Although LLMs are black boxes, we tried to explain them with a prompt approach.
@@ 9 Conclusion
This paper investigates how to generate more intuitively explainable and flexible reasoning task prompts. Our experimental results on the mathematical reasoning tasks significantly surpass those of prior zero-shot methods with the GPT-3.5 family. Also, on the common sense benchmark StrategyQA, HoT reached a high accuracy of 82.96% with GPT-3.5-turbo.